# MOWA: Multiple-in-One Image Warping Model

### v2

@@ Abstract
While recent image warping approaches achieved remarkable success on existing benchmarks, they still require training separate models for each specific task and cannot generalize well to different camera models or customized manipulations. To address diverse types of warping in practice, we propose a Multiple-in-One image WArping model (named MOWA) in this work. Specifically, we mitigate the difficulty of multi-task learning by disentangling the motion estimation at both the region level and pixel level. To further enable dynamic task-aware image warping, we introduce a lightweight point-based classifier that predicts the task type, serving as prompts to modulate the feature maps for more accurate estimation. To our knowledge, this is the first work that solves multiple practical warping tasks in one single model. Extensive experiments demonstrate that our MOWA, which is trained on six tasks for multiple-in-one image warping, outperforms state-of-the-art task-specific models across most tasks. Moreover, MOWA also exhibits promising potential to generalize into unseen scenes, as evidenced by cross-domain and zero-shot evaluations. The code and more visual results can be found on the project page: https://kangliao929.github.io/projects/mowa/.
@@ I Introduction
Image warping is essential in the field of computational imaging and computer vision, serving as the foundation for numerous applications, including image rectification [1, 2, 3, 4], image rectangling [5, 6, 7, 8], camera calibration [9, 10, 11, 12, 13], and 3D reconstructions [14, 15, 16, 17], etc. Enabling the manipulation of image data through processes such as scaling, rotation, and sheering allows for the seamless integration of diverse visual elements and the correction of optical imperfections. Moreover, image warping is indispensable in developing augmented reality (AR) and virtual reality (VR) applications [18, 19, 20], where it helps create immersive and realistic environments by accurately mapping textures and images onto 3D models.
Considering different inputs derived from different camera models or manipulation spaces, recent works integrate specific prior knowledge into their models to address the corresponding image warping tasks [6, 5, 21, 4, 22, 23, 24, 25, 26]. While these single-task approaches achieve significant progress, we found they suffer from two main limitations: (i) the lack of generalization and flexibility, which restricts their real-world applications since users are required to manually identify each input type and apply the appropriate single-task model. This process is time-consuming and challenging for non-professional users to judge. (ii) the substantial storage requirements for multiple task-specific models, which is impractical for some resource-limited platforms. Thus, it is crucial to develop a holistic framework capable of efficiently warping images from various camera models or manipulation spaces. Furthermore, many image warping tasks typically involve shared processes, such as motion estimation and content-aware perception. This indicates the possibility of developing a unified framework that incorporates these common image techniques.
@@
Given the fact that learning different structures of motion is non-trivial in one model and motion representations differ significantly across various tasks, we propose to disentangle the motion estimation at both the region level and pixel level. In this hierarchical architecture, we first estimate the control points of the thin-plate spline (TPS) model [27] with increasing refined numbers, in which the feature maps are progressively warped and rectified. Such a representation excels in approximating complex motions at a region level and enables high flexibility to various motion structures. Subsequently, the warped feature maps are fed into the decoder to predict the residual pixel-level displacement, which further improves the warping results for each task, especially in the image boundaries and details.
To enable MOWA to explicitly discriminate diverse input types, a lightweight point-based classifier is devised. Adding an extra classification network based on the image features is a straightforward solution but brings high computation and storage costs. Noticing the motion structures in different warping tasks possess their specific distribution, we leverage the middle product of the image warping framework, i.e., region-level control points, to directly learn the task type. It achieves comparable performance while allowing significant parameter reduction compared to the image-based classifier since only a few 2D points are needed. Then, the task label predicted by this point-based classifier is used to modulate the feature maps in the decoder, dynamically boosting task-aware image warpings using a prompt learning module. Prompts are a set of learnable parameters that encapsulate essential discriminative information about different types of input, which empower a single model to efficiently traverse and harness its vast parameter space to accommodate various warping requirements.
In the experiments, we trained MOWA on six typical tasks for multiple-in-one image warping. Experimental results demonstrated that it outperforms state-of-the-art (SotA) task-specific models in most tasks, even with comparable parameters of the network. In addition, MOWA allows the ability to generalize to unseen scenes, as evidenced by cross-domain evaluation (unfamiliar domains) and zero-shot evaluation (unseen tasks), indicating its robustness and adaptability across various scenarios. Our contributions can be summarized as follows:
@@
  • We propose to mitigate the difficulty of multi-task learning by decoupling the motion estimation in both the region level and pixel level. Moreover, a prompt learning module, guided by a lightweight point-based classifier, is designed to facilitate task-aware image warpings.
@@
The remainder of the paper is organized as follows: Section II reviews the related works. We then present the proposed multiple-in-one image warping framework in Section III. The experiments are provided in Section IV. Section V concludes this paper.
@@ III-A Problem Definition
In this study, we consider six representative and practical image types in the field of computational photography, including stitched images, rectified wide-angle images, unrolling shutter images, rotated images, fisheye images, and portrait photos, covering the mainstream practical image warping tasks. These types are further classified into two groups. The first four types (stitched images, rectified wide-angle images, unrolling shutter images, and rotated images) struggle with irregular boundaries as the original images are manipulated by some customized operations, such as image stitching, distortion correction, and rotation. Therefore, image rectangling is proposed to reshape these irregular boundaries while keeping the distribution of content unchanged. The last two types (fisheye images and portrait photos) show inherently geometric distortions imaged by special camera models, such as radial distortion in fisheye images and perspective distortion in portrait photos. Correcting these distortions is crucial to scene understanding and aesthetic appreciation.
Figure 2 shows the overall framework of the proposed MOWA. It takes the image and mask as input and estimates the TPS control points with increasingly refined numbers. In this region-level motion estimation, the feature maps are progressively warped and rectified. Subsequently, the warped features are fed into the decoder to predict a residual pixel-level motion. To enable task-aware and expandable capabilities, a lightweight point-based classifier and prompt learning module are designed. We elaborate on the details of each module in the proposed multiple-in-one image warping framework as follows.
@@ III-C Point-based Task Classifier
In addition to the task classification function, our point-based task classifier can further improve image warping performance. This improvement is due to the high-level guidance provided by the task classifier to the motion estimation module through gradient back-propagation. Figure 3 (right) depicts a typical example, in which the fisheye rectification result shows a less distorted shape, and the predicted control points of the stitched image are more tightly aligned to the image boundary. Compared to the vanilla baseline, the proposed point-based task classifier achieves an average improvement of $+0.35$ in PSNR metrics across various image warping tasks. More quantitative results are presented in Section IV-C.
@@ III-E Training Loss
where $\lambda_{Flow}$ and $\lambda_{Cls}$ are the hyper-parameters to balance different losses, both of them are empirically set to $0.1$.
In summary, the proposed multiple-in-one image warping framework bring the following benefits.
@@ IV-A Experimental Settings
Implementation Details. We train the proposed model using the Adam optimizer with the momentum terms of $(0.9,0.999)$ on 8 NVIDIA A100 GPUs. The learning rate starts with a linear warm-up in the first three epochs and then decays from $1e^{-4}$ to $1e^{-6}$ following a cosine schedule in the remaining epochs. The batch size is set as 64. The complete framework is trained with a fixed input size of $256\times 256$. At the first 10 epochs, we solely train and supervise the TPS prediction heads with the point-based task classifier. Afterwards, all modules are trained collectively. During inference, the proposed method supports image warping for any resolution by scaling the predicted TPS control points and residual flow.
Network Configuration. We design the image warping network based on the encoder-decoder architecture, enabling both region-level control point regression and pixel-level residual flow prediction. Specifically, the Transformer blocks with shifted windows [48, 49] are used in both the encoder and decoder except for the input projection layer and output projection layer. The basic dimension of channels is set to 32 and linearly increases along the layers in the encoder network, which is oppositely decreased to 2 in the decoder network. Moreover, the depths of each Transformer block are set to 2 and the head numbers of multi-head self-attention are $[1,2,4,8,16,16,8,4,2]$ along the whole layers. In TPS prediction heads, we adopt the convolution layers with different kernels to predict increasing numbers of control points, and the numbers are set to $10\times 10$, $12\times 12$, $14\times 14$, and $16\times 16$. The configuration details of these regression heads are listed in Table I. Such a design enables significant parameter reduction compared with the fully connected layers. For the lightweight point-based classifier, three 1D convolutional layers with channel dimensions of $256,256,512$ are used to extract the features of input and then three fully connected layers with unit numbers of $512,256,6$ are used to classify their task types.
@@ IV-B Comparison Results
Qualitative Comparison. As shown in Fig. 4, we visualize the comparison results of different methods on the testing datasets. These qualitative results demonstrate that our multiple-in-one method can handle various tasks, scenes, and resolutions well, compared with the SotA methods specially designed for each task. For example, for the rotated images, our method can rearrange the input to a rectangle one while keeping the original geometric layout reasonable. On the contrary, distorted buildings can be observed in the results of previous works [5, 6], in which the physical world rules such as the horizon are perturbed. For other rectangling tasks like the stitched image, unrolling shutter image, and rectified wide-angle image, our method shows a better visual appearance, especially in the image boundaries, allowing promising structural integrity among the comparison methods. In some challenging cases, such as the first and second rows in stitched images, the image boundaries are dramatically stretched, but our method can still warp the images to expected structures. One important reason is that MOWA learns the generalized warping strategy from different tasks since it can extract some common knowledge from them. In addition, our method mitigates the difficulty of motion estimation by disentangling it at both the region level and pixel level. Consequently, diverse structures of motions can be progressively approximated and the image details can be preserved. For the fisheye image and portrait photos, MOWA is capable of recovering the realistic geometric distribution from the inputs, despite the radial distortion or perspective distortion. Please refer to more visual comparison images, interactive warping visualizations, and dynamic warping results on the project page: https://kangliao929.github.io/projects/mowa/.
@@
Figure 5: Ablation study on the proposed motion estimation module. The predicted TPS control points are shown with the size of $10\times 10$, $12\times 12$, $14\times 14$, and $16\times 16$, from the left to right. The coarse results and final results are obtained by warping the input using the first control points and final flow (coupled with the last TPS points and residual flow), respectively.
@@ IV-C Ablation Study
Motion Estimation. It is challenging to estimate multiple motions in one model since the motion’s complexities and patterns significantly differ across various tasks. For this purpose, we proposed a flexible and hierarchical architecture to disentangle the motion estimation at the region level and pixel level. As shown in Fig. 5, better localization performance of the formed mesh can be achieved by increasing the number of TPS points. Besides, the pixel-level residual flow can provide a higher degree of freedom for the motion than only the region-level motion representation, improving the warping results, particularly in the image boundaries and details. Table III quantitatively demonstrates the effectiveness of the proposed hierarchical motion estimation module. We also found the upper bound occurs when continuously increasing the number of control points, e.g, the performance of four motion estimation heads to predict the size of $16\times 16$ TPS even worse than the size of $12\times 12$ TPS. This suggests the performance of multiple-in-one image warping would be limited without proper decoupling motion manner.
@@
Figure 7: Generalization evaluation of the proposed method: cross-domain evaluation (top) and zero-shot evaluation (bottom). The visualized flow and control points are both predicted by MOWA. In image retargeting results, the red dotted lines measure the stretch extent of the face or body by warping operations.
@@ IV-D Analysis on Multi-Task Learning and Prompt Learning
In the proposed method, a prompt learning module is designed to modulate the feature map with the soft task label, which helps to dynamically navigate its extensive parameter space to achieve task-aware image warping. By combining this module into MOWA, the averaged PSNR metric of all warping results gains $+0.21$dB improvements beyond the baseline.
@@
Figure 8: While our method is trained on the synthetic datasets with clean masks, it shows good robustness to the real-world data with noisy pseudo masks. The motion estimation of real-world images, e.g., the predicted control points, and warping flow, exhibit structurally reasonable distributions similar to those of the synthetic images.
From Figure 6 (a), we have the following three observations: (1) Different tasks show various levels of difficulty when training a multiple-in-one image warping model. For example, learning to warp the unrolling shutter image (task 3) is generally easier than other tasks, which shows the fastest convergence at the first 80 epochs. The reason is that the structures of unrolling shutter images are basically regular, where more than two boundaries are straight and do not need to be warped. On the contrary, learning to warp the stitched images (task 1), rectified wide-angle images (task 2), and portrait photos (task 6) is more challenging (slow convergences can be observed) since their boundaries or expected motions vary greatly in datasets. Especially in the portrait photo dataset, the numbers, shapes, and locations of the faces are quite diverse. (2) The multiple-in-one model tends to sacrifice the performance of some individual tasks to achieve the improvement of holistic warping performance. Particularly, the accuracy of warping the unrolling shutter image is dramatically reduced since the 80th epoch, but the performance of other tasks continues to improve. Such a trade-off among different tasks facilitates an overall improvement but leads to performance conflicts between some tasks. (3) The relationship of different tasks can be positive or negative. For example, the curves of the stitched image and rotated image show consistent trends during the MOWA’s training as they share the similar warping principle, i.e., rectangling the irregular image boundaries and keeping the content unchanged. Thus, meaningful interactions between these two tasks could happen if learning a multiple-in-one model. However, the curve of warping the fisheye image shows a converse trend to those of the stitched image and rotated image. The reason for this is that correcting the radial distortion present in fisheye images significantly alters the scene’s layout while leaving the image boundaries almost untouched. This approach contrasts with the foundational principles of tasks based on image rectangling. As a result, the imbalanced convergences can be noticed across these tasks with negative relationships.
@@
Figure 9: Failure cases of the proposed approach. It fails to accurately warp the input image with challenging image boundaries using a certain number of control points.

### v3

@@ Abstract
While recent image warping approaches achieved remarkable success on existing benchmarks, they still require training separate models for each specific task and cannot generalize well to different camera models or customized manipulations. To address diverse types of warping in practice, we propose a Multiple-in-One image WArping model (named MOWA) in this work. Specifically, we mitigate the difficulty of multi-task learning by disentangling the motion estimation at both the region level and pixel level. To further enable dynamic task-aware image warping, we introduce a lightweight point-based classifier that predicts the task type, serving as prompts to modulate the feature maps for more accurate estimation. To our knowledge, this is the first work that solves multiple practical warping tasks in one single model. Extensive experiments demonstrate that our MOWA, which is trained on six tasks for multiple-in-one single image warping, outperforms state-of-the-art task-specific models across most tasks. Moreover, MOWA also exhibits promising potential to generalize into unseen scenes, as evidenced by cross-domain and zero-shot evaluations. The code and more visual results can be found on the project page: https://kangliao929.github.io/projects/mowa/.
@@ I Introduction
Image warping is essential in the field of computational imaging and computer vision, serving as the foundation for numerous applications, including image rectification [1, 2, 3, 4], image rectangling [5, 6, 7, 8], camera calibration [9, 10, 11, 12, 13], and 3D reconstructions [14, 15, 16, 17], etc. Enabling the manipulation of image data through processes such as scaling, rotation, and shearing allows for the seamless integration of diverse visual elements and the correction of optical imperfections. Moreover, image warping is indispensable in developing augmented reality (AR) and virtual reality (VR) applications [18, 19, 20], where it helps create immersive and realistic environments by accurately mapping textures and images onto 3D models.
Considering different inputs derived from different camera models or manipulation spaces, recent works integrate specific prior knowledge into their models to address the corresponding image warping tasks [6, 5, 21, 4, 22, 23, 24, 25, 26]. While these single-task approaches achieve significant progress, we found they suffer from two main limitations: (i) the lack of generalization and flexibility, which restricts their real-world applications since users are required to manually identify each input type and apply the appropriate single-task model. This process is time-consuming and challenging for non-professional users to judge; (ii) the substantial storage requirements for multiple task-specific models, which are impractical for some resource-limited platforms. Thus, it is crucial to develop a holistic framework capable of efficiently warping images from various camera models or manipulation spaces. Furthermore, many image warping tasks typically involve shared processes, such as motion estimation and content-aware perception. This indicates the possibility of developing a unified framework that incorporates these common image techniques.
@@
Given the fact that learning different structures of motion is non-trivial in one model and motion representations differ significantly across various tasks, we propose to disentangle the motion estimation at both the region level and pixel level. In this hierarchical architecture, we first estimate the control points of the thin-plate spline (TPS) model [27] with increasing refined numbers, in which the feature maps are progressively warped and rectified. Such a representation excels in approximating complex motions at region level  [^1] The region specifically denotes each grid formed by TPS control points rather than the segmented parts in some vision tasks. and enables high flexibility to various motion structures. Subsequently, the warped feature maps are fed into the decoder to predict the residual pixel-level displacement, which further improves the warping results for each task, especially in the image boundaries and details.
To enable MOWA to explicitly discriminate diverse input types, a lightweight point-based classifier is devised. Adding an extra classification network based on the image features is a straightforward solution but brings high computation and storage costs. Noticing the motion structures in different tasks possess their specific distribution, we leverage the middle product of the image warping framework, i.e., region-level control points, to directly learn the task type. It achieves comparable performance while allowing significant parameter reduction compared to the image-based classifier since only a few 2D points are needed. Then, the task label predicted by this point-based classifier is used to modulate the feature maps in the decoder, dynamically boosting task-aware image warpings using a prompt learning module. Prompts are a set of learnable parameters that encapsulate essential discriminative information about different types of input, which empower a single model to efficiently traverse and harness its vast parameter space to accommodate various warping requirements.
In the experiments, we trained MOWA on six typical tasks for multiple-in-one single image warping. Experimental results demonstrated that it outperforms state-of-the-art (SotA) task-specific models in most tasks, even with comparable network parameters. In addition, MOWA allows the ability to generalize to unseen scenes, as evidenced by cross-domain evaluation (unfamiliar domains) and zero-shot evaluation (unseen tasks), indicating its robustness and adaptability across various scenarios. Our contributions can be summarized as follows:
@@
  • We propose to mitigate the difficulty of multi-task learning by decoupling the motion estimation at both the region level and pixel level. Moreover, a prompt learning module, guided by a lightweight point-based classifier, is designed to facilitate task-aware image warpings.
@@
The remainder of the paper is organized as follows: Section II reviews the related works. We then present the proposed MOWA in Section III. The experiments are provided in Section IV. Section V concludes this paper.
@@ III-A Problem Definition
In this study, we consider six representative and practical image types in the field of computational photography, including stitched images, rectified wide-angle images, unrolling shutter images, rotated images, fisheye images, and portrait photos, covering the mainstream practical image warping tasks. These types are further classified into two groups. The first four types (stitched images, rectified wide-angle images, unrolling shutter images, and rotated images) struggle with irregular boundaries as the original images are manipulated by some customized operations, such as image stitching, distortion correction, and rotation. Therefore, image rectangling is proposed to reshape these irregular boundaries while keeping the distribution of content unchanged. The last two types (fisheye images and portrait photos) show inherently geometric distortions imaged by special camera models, such as radial distortion in fisheye images and multiple distortions (both radial distortion and perspective distortion) in portrait photos. Correcting these distortions is crucial to scene understanding and aesthetic appreciation.
Figure 2 shows the overall framework of the proposed MOWA. It takes the image and mask as input and estimates the TPS control points with increasingly refined numbers. In this region-level motion estimation, the feature maps are progressively warped and rectified. Subsequently, the warped features are fed into the decoder to predict a residual pixel-level motion. To enable task-aware and expandable capabilities, a lightweight point-based classifier and prompt learning module are designed. We elaborate on the details of each module in MOWA as follows.
@@ III-C Point-based Task Classifier
In addition to the task classification function, our point-based task classifier can further improve image warping performance. This improvement is due to the high-level guidance provided by the task classifier to the motion estimation module through gradient back-propagation. Figure 3 (right) depicts a typical example, in which the fisheye rectification result shows a less distorted shape, and the predicted control points of the stitched image are more tightly aligned to the image boundary. Compared to the vanilla baseline, the proposed point-based task classifier achieves an average improvement of $+0.35$dB in PSNR metrics across various image warping tasks. More quantitative results are presented in Section IV-C.
@@ III-E Training Loss
where $\lambda_{Flow}$ and $\lambda_{Cls}$ are the hyperparameters to balance different losses, both of them are empirically set to $0.1$.
In summary, the proposed multiple-in-one image warping framework brings the following benefits.
@@ IV-A Experimental Settings
Implementation Details. We train the proposed model using the AdamW optimizer with the momentum terms of $(0.9,0.999)$ on 8 NVIDIA A100 GPUs. The learning rate starts with a linear warm-up in the first three epochs and then decays from $1e^{-4}$ to $1e^{-6}$ following a cosine schedule in the remaining epochs. The batch size is set as 64. The complete framework is trained with a fixed input size of $256\times 256$. At the first 10 epochs, we solely train and supervise the TPS prediction heads with the point-based task classifier. Afterwards, all modules are trained collectively. During inference, the proposed method supports image warping for any resolution by scaling the predicted TPS control points and residual flow.
Network Configuration. We design the image warping network based on the encoder-decoder architecture, enabling both region-level control point regression and pixel-level residual flow prediction. Specifically, the Transformer blocks with shifted windows [48, 49] are used in both the encoder and decoder except for the input projection layer and output projection layer. The basic dimension of channels is set to 32 and linearly increases along the layers in the encoder network, which is oppositely decreased to 2 in the decoder network. Moreover, the depths of each Transformer block are set to 2 and the head numbers of multi-head self-attention are $[1,2,4,8,16,16,8,4,2]$ along the whole layers. In TPS prediction heads, we adopt the convolution layers with different kernels to predict increasing numbers of control points, and the numbers are set to $10\times 10$, $12\times 12$, $14\times 14$, and $16\times 16$. The configuration details of these regression heads are listed in Table I. Such a design enables significant parameter reduction compared with the fully connected layers. For the lightweight point-based classifier, three 1D convolutional layers with channel dimensions of $256,256,512$ are used to extract the features of input, and then three fully connected layers with unit numbers of $512,256,6$ are used to classify their task types.
@@ IV-B Comparison Results
Qualitative Comparison. As shown in Fig. 4, we visualize the comparison results of different methods on the testing datasets. These qualitative results demonstrate that our multiple-in-one method can handle various tasks, scenes, and resolutions well, compared with the SotA methods specially designed for each task. For example, for the rotated images, our method can rearrange the input to a rectangle one while keeping the original geometric layout reasonable. On the contrary, distorted buildings can be observed in the results of previous works [5, 6], in which the physical world rules such as the horizon are perturbed. For other rectangling tasks like the stitched image, unrolling shutter image, and rectified wide-angle image, our method shows a better visual appearance, especially in the image boundaries, allowing promising structural integrity among the comparison methods. In some challenging cases, such as the first and second rows in stitched images, the image boundaries are dramatically stretched, but our method can still warp the images to the expected structures. One important reason is that MOWA learns the generalized warping strategy from different tasks since it can extract some common knowledge from them. In addition, our method mitigates the difficulty of motion estimation by disentangling it at both the region level and pixel level. Consequently, diverse structures of motions can be progressively approximated and the image details can be preserved. For the fisheye image and portrait photos, MOWA is capable of recovering the realistic distribution from the inputs with distorted objects or irregular boundaries, despite the radial distortion or perspective distortion. Please refer to more visual comparison images, interactive warping visualizations, and dynamic warping results on the project page: https://kangliao929.github.io/projects/mowa/.
@@
Figure 5: Ablation study on the proposed motion estimation module. The predicted TPS control points are shown with the size of $10\times 10$, $12\times 12$, $14\times 14$, and $16\times 16$, from left to right. The coarse results and final results are obtained by warping the input using the first control points and final flow (coupled with the last TPS points and residual flow), respectively.
@@ IV-C Ablation Study
Motion Estimation. It is challenging to estimate multiple motions in one model since the motion’s complexities and patterns significantly differ across various tasks. For this purpose, we proposed a flexible and hierarchical architecture to disentangle the motion estimation at the region level and pixel level. As shown in Fig. 5, better localization performance of the formed mesh can be achieved by increasing the number of TPS points. Besides, the pixel-level residual flow can provide a higher degree of freedom for the motion than only the region-level motion representation, improving the warping results, particularly in the image boundaries and details. Table III quantitatively demonstrates the effectiveness of the proposed hierarchical motion estimation module. We also found the upper bound occurs when continuously increasing the number of control points, e.g, the performance of four motion estimation heads to predict the size of $16\times 16$ TPS is even worse than the size of $12\times 12$ TPS. This suggests the performance of multiple-in-one image warping would be limited without proper decoupling of motion manner.
@@
TABLE V: The ablation study of various combinations involved with different warping tasks. We choose the first combination of any two tasks (‘Task 1-Stitched Image’ with others) as the comparison baseline. The performance gains of PSNR/SSIM are reported for each combination.
Task 1 Task 2 Task 3 Task 4 Task 5
Task 1 - +0.0/0.0 +0.27/0.0210 -0.12/0.0077 +0.22/0.0198
Task 2 +$0$ +0.32/0.0190 +0.22/0.0129 +0.25/0.0151
Task 3 +0.0/0.0 +0.11/0.0077 - +0.08/0.034 +0.19/0.0087
Task 4 +0.0/0.0 +0.26/0.0129 +0.55/0.0276 - +0.47/0.0218
Task 5 +0.0/0.0 +0.98/0.0377 +0.20/0.0097 +0.79/0.0285 -
@@
Figure 7: Generalization evaluation of the proposed method: cross-domain evaluation (top) and zero-shot evaluation (bottom). The visualized flow and control points are both predicted by MOWA. The image resolution is marked for each real-world image. In image retargeting results, the red dotted lines measure the stretch extent of the face or body by warping operations.
@@ IV-D Analysis on Multi-Task Learning and Prompt Learning
In the proposed method, a prompt learning module is designed to modulate the feature map with the learned task label, which helps to dynamically navigate its extensive parameter space to achieve task-aware image warping. By combining this module into MOWA, the averaged PSNR metric of all warping results gains $+0.21$dB improvements beyond the baseline.
@@
Figure 8: While our method is trained on the synthetic datasets with clean masks, it shows good robustness to the real-world data with noisy pseudo masks. The motion estimations of real-world images exhibit structurally reasonable distributions similar to those of the synthetic images.
Figure 9: Extended multi-view applications of the proposed method: image alignment, image stitching, and stereo image rectification. In the fusion results of image alignment, the non-overlapping regions of two views are marked in orange. The linearly average fusion algorithm is applied for the image stitching. In the stereo image rectification, the red dotted lines are drawn to highlight the vertical alignment. The stereo matching results are inferred by HITNet [51].
From Figure 6 (a), we have the following three observations: (1) Different tasks show various levels of difficulty when training a multiple-in-one image warping model. For example, learning to warp the unrolling shutter image (task 3) is generally easier than other tasks, which shows the fastest convergence at the first 80 epochs. The reason is that the structures of unrolling shutter images are basically regular, where more than two boundaries are straight and do not need to be warped. On the contrary, learning to warp the stitched images (task 1), rectified wide-angle images (task 2), and portrait photos (task 6) is more challenging (slow convergences can be observed) since their boundaries or expected motions vary greatly in datasets. Especially in the portrait photo dataset, the numbers, shapes, and locations of the faces are quite diverse. (2) The multiple-in-one model tends to sacrifice the performance of some individual tasks to achieve the improvement of holistic warping performance. Particularly, the accuracy of warping the unrolling shutter image is dramatically reduced since the 80th epoch, but the performance of other tasks continues to improve. Such a trade-off among different tasks facilitates an overall improvement but leads to performance conflicts between some tasks. (3) The relationship of different tasks can be positive or negative. For example, the curves of the stitched image and rotated image show consistent trends during the MOWA’s training as they share the similar warping principle, i.e., rectangling the irregular image boundaries and keeping the content unchanged. Thus, meaningful interactions between these two tasks could happen if learning a multiple-in-one model. However, the curve of warping the fisheye image shows a converse trend to those of the stitched image and rotated image. The reason for this is that correcting the radial distortion present in fisheye images significantly alters the scene’s layout while leaving the image boundaries almost untouched. This approach contrasts with the foundational principles of tasks based on image rectangling. As a result, the imbalanced convergences can be noticed across these tasks with negative relationships. Moreover, the ablation study of exploring various combinations involved with different warping tasks is shown in Table V. Specifically, we choose the first combination of any two tasks (‘Task 1’ with others) as the comparison baseline and report the performance gains of PSNR/SSIM for each subsequent combination.
@@
Figure 10: Failure cases of the proposed method. It fails to accurately warp the input image with challenging image boundaries using a certain number of control points.